# HAFixAgent: History-Aware Automated Program Repair Agent

### v2

@@ Abstract.
Automated program repair (APR) has recently shifted toward large language models and agent-based systems, yet most systems rely on local snapshot context, overlooking repository history. Prior work shows that repository history helps repair single-line bugs, since the last commit touching the buggy line is often the bug-introducing one. In this paper, we investigate whether repository history can also improve agentic APR systems at scale, especially for complex multi-hunk bugs. We present HAFixAgent, a History-Aware Bug-Fixing Agent that injects blame-derived repository heuristics into its repair loop. A preliminary study of all 854 real-world bugs from Defects4J motivates our design, showing that bug-relevant history is both widely available and highly concentrated. Empirical comparison of HAFixAgent with two state-of-the-art baselines shows: (1) Effectiveness: HAFixAgent significantly improves over the agent-based baseline (by 212.3%) and the multi-hunk baseline (by 29.9%). (2) Efficiency: history does not significantly increase agent steps and keeps token costs comparable, with notably lower median costs for complex multi-file-multi-hunk bugs. (3) Practicality: combining different historical heuristics repairs more bugs, offering a clear cost-benefit trade-off. HAFixAgent offers a practical recipe for history-aware agentic APR: ground the agent in version control history, prioritize diff-based historical context, and integrate complementary heuristics when needed.
@@ 1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (Boulder2013ReverseDebug; zhang2023survey; weiss2007long). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (goues2019apr). APR has evolved from heuristic-based (GenProg; le2016history; wen2018context), constraint-based (le2017s3; mechtaev2016angelix; long2015staged), and template-based techniques (liu2019tbar; martinez2016astor; liu2019avatar) to learning-based methods (chen2019sequencer; jiang2021cure; li2020dlfix; lutellier2020coconut; zhu2021syntax; zhu2023tare; ye2024iter) and, most recently, large language model-based approaches (zhang2024systematic; jiang2023impact; hossain2024deep; xia2023automated; xia2024automated; xia2023plastic; fan2023automated; yin2024thinkrepair; huang2025template; li2024hybrid; shi2025hafix) and agent-based approaches (yang2024sweagent; bouzenia2024repairagent; mu2025experepair; meng2024empirical; lee2024unified; zhang2024autocoderover; chen2025swe; wang2024openhands). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
@@
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, potentially non-local edits across different parts of the codebase. Recent characterizations, such as the work by nashid2025characterizing, highlight that these bugs are not only common but also significantly harder for current LLM-based approaches compared to simpler single-hunk fixes. Overcoming this complexity barrier is crucial for APR approaches to be broadly applicable in real-world software development.
@@
To address such complex bugs, recent advances in APR have pursued two promising yet separate directions. First, agent-based approaches have emerged as the state-of-the-art (bouzenia2024repairagent; yang2024sweagent; wang2024openhands). Agent-based approaches like RepairAgent (bouzenia2024repairagent) leverage LLMs to plan, execute actions via software tools (e.g., code search, editing, testing), and react to iterative feedback, enabling them to tackle more complex reasoning compared to simple prompting. However, these agents typically operate only on the current code snapshot and immediate test results, lacking deeper historical context. Second, drawing from decades of mining software repositories research (sliwerski2005changes; hassan2006mining; adams2010identifying), recent APR techniques started to recognize the value of historical code context. Our prior work (shi2025hafix) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits (the last change touching a buggy line) can improve LLM repair performance for single-line bugs using simple prompting. However, this approach lacked the sophisticated reasoning and tool interaction capabilities of agents, limiting its potential for complex, multi-location repairs.
@@
We hypothesize that combining the iterative, tool-using capabilities of agentic workflows with the rich, contextual clues found in repository history can significantly enhance APR, particularly for complex multi-hunk bugs. History provides information on relevant past changes, developer rationale, and co-evolution patterns, while the agent leverages tools to explore, validate, and integrate this historical context dynamically during the repair process. Yet, it is unclear whether this approach scales. Complex multi-hunk bugs could have many different blame commits, potentially overwhelming the agent with noisy and irrelevant signals rather than providing a focused context.
@@
We design HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by our prior HAFix work (shi2025hafix)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in the Defects4J (just2014defects4j) dataset, a widely used benchmark for evaluating APR techniques. In this study, we address the following research questions (RQs):
@@
  • RQ0. To What Extent Is Historical Information Available in Real-World Bugs? Before integrating repository history, we have to first establish if relevant historical information is broadly available and scalable for use in APR. Specifically, we analyze the Defects4J dataset to measure the availability and concentration of historical information for bugs. Our findings confirm that 71.1% of bugs have available historical information and most map to a single unique blame commit. This result directly supports the feasibility of using a simple, single-commit heuristic for HAFixAgent, even for complex bugs.
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent to recent APR systems and also quantify the extent to which any improvements are attributable to the introduction of historical context as opposed to the adoption of an agent architecture. For this, we evaluate HAFixAgent’s effectiveness against two state-of-the-art baselines on the Defects4J dataset: one for common bugs (bouzenia2024repairagent) and another specializing in multi-hunk repairs (nashid2025characterizing). The results show that HAFixAgent repairs 212.3% more bugs than the baseline for common bugs and outperforms the multi-hunk baseline by 29.9%.
  • RQ2. What Are the Cost and Efficiency of HAFixAgent? Historical grounding should not inflate agent loops or token usage. We track token costs of HAFixAgent under different configurations and compare the costs. We found that integrating history into the agent repair loop is cost-efficient overall, without notable overhead.
@@
  • A practical and automated method for constructing historical context, featuring three distinct blame-derived historical heuristics and a robust fallback strategy designed for any agentic APR system.
  • A large-scale empirical evaluation on all 854 Defects4J bugs, providing strong evidence that history is (1) broadly available (71.1% blameable) and highly concentrated (with 70.7% of bugs mapping to a single commit), and (2) significantly improves agent effectiveness and efficiency.
We will release the complete replication package of code and the evaluation script after the revision is completed. In the interim, please contact us by email to request access.
@@
Section 2 provides background information. Section 3 presents the preliminary study on history availability and distribution. Section 4 describes our architecture design of HAFixAgent. Section 5 presents our case study design. Section 6 presents the results. Section 7 discusses the possible future work. Section 8 outlines threats to the validity of our study. Section 9 reviews related work. Section 10 concludes this paper.
@@ 2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (prenner2024out; sintaha2023katana). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. hassan2006mining demonstrated the broad utility of mining repositories to assist developers. More specifically, sliwerski2005changes showed that analyzing the change-inducing commit can help locate faults. adams2010identifying used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
@@
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, le2016history pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, kamei2012large reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
@@
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (mu2025experepair) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (chen2025swe) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
@@
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (sliwerski2005changes; kim2006automatic). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (shi2025hafix), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
@@ 2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (openai2025gpt5) and DeepSeek-V3 (liu2024deepseek), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (wang2024survey). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (yao2023react). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (hassan2024rethinking). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (hassan2025agentic; li2025rise). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (roychoudhury2025agentic).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (yang2024sweagent) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (wang2024openhands) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (bouzenia2024repairagent) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (zhang2024autocoderover) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
@@ 3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (prenner2024out). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (prenner2024out; haque2025towards; fakhoury2024nl2fix; chen2024large; xia2023automated; li2024hybrid), only recently APR techniques were proposed that leverage code history (shi2025hafix; ehsani2025bug). Inspired by the domain of mining software repositories (sliwerski2005changes; hassan2006mining; kamei2012large), our prior work on single-line APR leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as the latter only comprise one buggy line.
@@ 3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs, we choose the latest version of the Defects4J dataset (v3.0.1) (just2014defects4j), a standard and widely-used benchmark for evaluating recent APR techniques (xia2022less; jiang2023impact; hossain2024deep; lutellier2020coconut; jiang2021cure; ye2024iter; bouzenia2024repairagent). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). Analyzing the entire dataset allows us to evaluate the general history distribution on different projects and bugs, without restricting to small or specific bug categories.
Consistent with common practices in the APR field (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug), we assume perfect fault localization. Specifically, we use the developer-written patch provided by Defects4J to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, our approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.

### v3

# HAFixAgent: History-Aware Program Repair Agent

@@ Abstract.
Automated program repair (APR) has recently shifted toward large language models and agent-based systems, yet most systems rely on local snapshot context, overlooking repository history. Prior work shows that repository history helps repair single-line bugs, since the last commit touching the buggy line is often the bug-introducing one. In this paper, we investigate whether repository history can also improve agentic APR systems at scale, especially for complex multi-hunk bugs. We present HAFixAgent, a History-Aware Bug-Fixing Agent that injects blame-derived repository heuristics into its repair loop. A preliminary study on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs motivates our design, showing that bug-relevant history is widely available across both benchmarks. Using the same LLM (DeepSeek-V3.2-Exp) for all experiments, including replicated baselines, we show: (1) Effectiveness: HAFixAgent outperforms RepairAgent (+56.6%) and BIRCH-feedback (+47.1%) on Defects4J. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, especially on single-file multi-hunk (SFMH) bugs. (2) Robustness: under noisy fault localization (+1/+3/+5 line shifts), history provides increasing resilience, maintaining 40 to 56% success on SFMH bugs where the non-history baseline collapses to 0%. (3) Efficiency: history does not significantly increase agent steps or token costs on either benchmark.
@@ 1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (BOULDER, 2013; Zhang et al., 2024a; Weiß et al., 2007). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (Le Goues et al., 2019). APR has evolved from heuristic-based (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018), constraint-based (Le et al., 2017; Mechtaev et al., 2016; Long and Rinard, 2015), and template-based techniques (Liu et al., 2019b; Martinez and Monperrus, 2016; Liu et al., 2019a) to learning-based methods (Chen et al., 2021b; Jiang et al., 2021; Li et al., 2020; Lutellier et al., 2020; Zhu et al., 2021, 2023; Ye and Monperrus, 2024) and, most recently, large language model-based approaches (Zhang et al., 2024b; Jiang et al., 2023; Hossain et al., 2024; Xia et al., 2023b; Xia and Zhang, 2024; Xia et al., 2023a; Fan et al., 2023; Yin et al., 2024; Huang et al., 2025; Li et al., 2025a; Shi et al., 2025) and agent-based approaches (Yang et al., 2024; Bouzenia et al., 2025; Mu et al., 2025; Meng et al., 2024; Lee et al., 2024; Zhang et al., 2024c; Chen et al., 2025; Wang et al., 2025). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
@@
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, non-local edits. As Nashid et al. (2025) highlight, these bugs are common yet significantly harder for current LLM-based approaches.
@@
To address such complex bugs, recent APR advances have pursued two promising yet separate directions. First, agent-based approaches (Bouzenia et al., 2025; Yang et al., 2024; Wang et al., 2025) leverage LLMs with iterative tool use (code search, editing, testing) to tackle complex reasoning, but typically operate only on the current code snapshot without deeper historical context. Second, drawing from decades of mining software repositories (MSR) research (Sliwerski et al., 2005; Hassan, 2006; Adams et al., 2010), prior work (Shi et al., 2025) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits can improve LLM repair for single-line bugs, but lacked the iterative reasoning capabilities of agents.
@@
We hypothesize that combining agentic workflows with repository history can significantly enhance APR, particularly for complex multi-hunk bugs. History provides information on relevant past changes, developer rationale, and co-evolution patterns, while the agent leverages tools to explore and validate repairs. Yet, it is unclear whether this scales, as complex bugs could have many blame commits, potentially overwhelming the agent with noisy signals.
@@
This paper presents HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by HAFix (Shi et al., 2025)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in Defects4J (Just et al., 2014) (Java) and all 501 bugs in BugsInPy (Widyasari et al., 2020) (Python), two widely used benchmarks for evaluating APR techniques. In this study, we address the following research questions (RQs):
@@
  • RQ0. To What Extent Is Historical Information Available in Real-World Bugs? Before integrating repository history, we first establish if relevant historical information is broadly available and scalable for use in APR. We analyze both Defects4J and BugsInPy to measure the availability and concentration of historical information. Our findings confirm that 71.1% (Defects4J) and 87.4% (BugsInPy) of bugs have available historical information. In Defects4J, 99.5% of blameable bugs map to a single blame commit, while BugsInPy shows more scattered blame (58.4% single-commit).
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent against two state-of-the-art baselines replicated on the same LLM (DeepSeek-V3.2-Exp) for a fair comparison, and perform a controlled ablation (history vs. non-history) on both Defects4J and BugsInPy. HAFixAgent outperforms RepairAgent by +56.6% and BIRCH-feedback by +47.1%. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, with the three history heuristics solving complementary sets of bugs.
  • RQ2. How Efficient and Robust Is History-Aware Repair? A practical concern is whether history-aware repair remains effective under imperfect fault localization and whether it inflates agent cost. We evaluate robustness by applying line-shift noise (+1/+3/+5) to fault localization on a stratified sample of 100 Defects4J bugs, and analyze cost/efficiency across both benchmarks. We find that history context provides increasing resilience under noisy FL, and does not significantly inflate cost.
@@
  • A practical and automated method for constructing historical context, featuring three distinct blame-derived historical heuristics and a robust fallback strategy for any agentic APR system.
  • A large-scale empirical evaluation on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs, with same-LLM baseline replication, providing strong evidence that history (1) is broadly available across languages, (2) significantly improves agent effectiveness, (3) provides increasing resilience under noisy fault localization, and (4) does not inflate cost.
@@
Section 2 provides background information. Section 3 presents the preliminary study on history availability and distribution. Section 4 describes our architecture design of HAFixAgent. Section 5 presents our case study design. Section 6 presents the results. Section 7 discusses the implications. Section 8 outlines threats to the validity of our study. Section 9 reviews related work. Section 10 concludes this paper.
@@ 2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (Prenner and Robbes, 2024; Sintaha et al., 2023). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. Hassan (2006) demonstrated the broad utility of mining repositories to assist developers. More specifically, Sliwerski et al. (2005) showed that analyzing the change-inducing commit can help locate faults. Adams et al. (2010) used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
@@
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, Le et al. (2016) pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, Kamei et al. (2013) reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
@@
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (Mu et al., 2025) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (Chen et al., 2025) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
@@
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (Sliwerski et al., 2005; Kim et al., 2006). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (Shi et al., 2025), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
@@ 2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (OpenAI, 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2024), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (Wang et al., 2024). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (Yao et al., 2023). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (Hassan et al., 2024). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (Hassan et al., 2025; Li et al., 2025b). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (Roychoudhury, 2025).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (Yang et al., 2024) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (Wang et al., 2025) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (Bouzenia et al., 2025) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (Zhang et al., 2024c) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
@@ 3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (Prenner and Robbes, 2024). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (Prenner and Robbes, 2024; Haque et al., 2025; Fakhoury et al., 2024; Chen et al., 2024; Xia et al., 2023b; Li et al., 2025a), only recently APR techniques were proposed that leverage code history (Shi et al., 2025; Ehsani et al., 2025). Inspired by the domain of mining software repositories (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013), prior work on single-line APR (Shi et al., 2025) leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as they only comprise one buggy line.
@@ 3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs across languages, we choose two established benchmarks. Defects4J (v3.0.1) (Just et al., 2014) is a standard and widely-used benchmark for evaluating recent APR techniques (Xia and Zhang, 2022; Jiang et al., 2023; Hossain et al., 2024; Lutellier et al., 2020; Jiang et al., 2021; Ye and Monperrus, 2024; Bouzenia et al., 2025). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). BugsInPy (Widyasari et al., 2020) is a benchmark of 501 Python bugs collected from 17 popular open-source projects, including data processing (pandas), web frameworks (fastapi, tornado), and utility libraries (tqdm, thefuck). BugsInPy complements Defects4J by covering a dynamically typed language with different development practices and commit patterns.
Consistent with common practices in the APR field (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we use the developer-written patch provided by each benchmark to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, the approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.